OpenStudy (anonymous):

What is the volume of the box if it is scaled down by a factor of ½? The volume of the box is 473. How do I solve this?

OpenStudy (anonymous):

if you know the original Volume is 473, how do you scale it down by 1/2?

OpenStudy (anonymous):

Do I divied bye 1/2 for the length width and height?

OpenStudy (bradely):

Volume = (1/2)^3 (Original volume) Volume = (1/2)^3 (473) Source: http://www.mathskey.com/question2answer/